100 mL of buffer solution that is 0.15M HA (Ka= 6.8x10-5) and 0.20M NaA is mixed with 18.2mL of 0.35M NaOH. What is the pH of the resulting solution?
Answer #1

pKa of acid = -log Ka = -log 6.8x10-5 =4.1675

The buffer is

HA + NaOH ----------------------> NaA + H2O

100x0.15=15 0 100x 0.20=20 - ---- initial mmoles

--------- 18.2 x0.35=6.37 ---------- ------- change

15-6.37=8.63 0 20+6.37=26.37 ------- equilibrium mmoles

Thus the solution is still a buffer and its pH is calculated using Hendersen equation as

pH = pKa + log [conjugate base]/[acid]

=4.1675 + log 26.37/8.63 =4.6526

The pH of the resulting solution = 4.6526